Write the Solubility Product Constant (Ksp) expression for the following solution at 275 oC.

MgF2(s)Mg2+(aq) + 2 F-(aq)
1
Expert's answer
2019-08-20T08:23:50-0400

"K_{sp}=[Mg^{2+}][F^-]^2"

At "25^0C" "K_{sp}=5.6\times10^{-11}"


At "275^oC" we should use Van't Hoff equation

"lnK = (\frac{-\Delta H^o}{R})\times \frac{1}{T} + \frac{\Delta S^0}{R}"